### v1

# Exploring Radio Emissions from Confirmed Exoplanets Using SKA

Abstract
Currently, our understanding of magnetic fields in exoplanets remains limited compared to those within our solar system. Planets with magnetic fields emit radio signals primarily due to the Electron Cyclotron Maser Instability mechanism. In this study, we explore the feasibility of detecting radio emissions from exoplanets using the Square Kilometre Array (SKA) radio telescope. Utilizing data from the NASA Exoplanet Archive, we compile information on confirmed exoplanets and estimate their radio emissions using the RBL model. Our analysis reveals that three exoplanets— Qatar-4 b, TOI-1278 b, and WASP-173 A b —exhibit detectable radio signals suitable for observation with the SKA telescope.
journal: Frontiers
1 Introduction
Observing and measuring the magnetic fields of exoplanets holds paramount importance in contemporary astrophysical research. The magnetic field of a planet serves as a fundamental aspect of its physical makeup, providing crucial insights into its internal structure and dynamical processes. Understanding exoplanetary magnetic fields allows for the investigation of planetary interiors, shedding light on the presence of metallic cores and the mechanisms driving magnetic field generation, such as dynamo processes. Additionally, the interaction between a planet’s magnetic field and the surrounding stellar wind profoundly influences its atmospheric dynamics, including the retention and loss of atmospheric gases. This interplay is pivotal in determining a planet’s habitability and its potential to support life as we know it. Furthermore, studying exoplanetary magnetic fields offers a glimpse into the diverse array of planetary systems beyond our own, enriching our understanding of planetary formation and evolution across different stellar environments and orbital configurations.
However, the task of observing and measuring the magnetic field of exoplanets presents a significant challenge. The magnetospheres of several solar system bodies, such as Earth, Jupiter, Saturn, Uranus, and Neptune, manifest distinct patterns of nonthermal continuum radiation emission due to their interaction with the solar wind (Zarka, 1998). The electron cyclotron maser instability (ECMI) mechanism is believed to be the primary source of non-thermal radio emission in planetary magnetospheres, where energetic electrons emit coherent radiation along magnetic field lines (Melrose & Dulk, 1982; Dulk, 1985). Observations of such emissions provide valuable insights into the strength and configuration of exoplanetary magnetic fields. However, despite extensive efforts, direct detection of radio emissions from exoplanets has proven challenging due to the low sensitivity of ground-based radio telescopes (Murphy et al., 2015; Lazio et al., 2018; Turner et al., 2021). Recent advancements, particularly in the development of next-generation instruments like the Square Kilometre Array (SKA), offer promising opportunities for more sensitive and comprehensive observations of exoplanetary radio emissions. By detecting and characterizing these emissions, we can gain deeper insights into the magnetic properties, atmospheric dynamics, and potentially the presence of exomoons in exoplanetary systems.
To date, the confirmed count of exoplanets has reached  5300. In this study, we conduct an assessment of the non-thermal radio emissions originating from these exoplanets, utilizing the RBL model. Subsequently, we identify potential candidates for detection using the Square Kilometre Array (SKA) telescope.
2 Methodology
2.1 Estimation of Exoplanets Radio Emission
Observations of magnetized planets within our Solar System have revealed a consistent relationship between the power of auroral emissions and the energy flux of the incoming stellar wind (Zarka et al., 2001; Zarka, 2004, 2007; Zarka et al., 2018). This phenomenon extends to various planetary bodies, including those with satellite-Jupiter radio emissions (Noyola et al., 2014, 2016). Specifically, studies have shown that approximately 1% of the energy carried by electrons in these systems manifests as radio waves (Zarka et al., 2018). This correlation between auroral emissions and stellar wind energy flux, along with the strength of the planetary magnetic field, is encapsulated by the RBL model (Zarka et al., 2001; Zarka, 2004, 2007). Mathematically, this model relates the emitted frequency to the magnetic field strength, as expressed by
$\nu=\frac{eB}{2\pi m}~{},$ (1)
where $\nu$ is the emitted frequency, $e$ is the charge of the emitting particles (in this case, electrons), $B$ is the magnetic field strength and $m$ is the mass of the emitting electrons.
Similarly, the radio power from an exoplanet is related to the incident power of the stellar winds as
$P_{rad}=4\times 10^{18}erg~{}s^{-1}\big{(}\frac{{\dot{M}_{ion}}}{10^{-14}M_{\odot}~{}yr^{-1}}\big{)}^{0.8}\big{(}\frac{v_{\infty}}{400~{}km~{}s^{-1}}\big{)}^{2}\big{(}\frac{a}{5~{}au}\big{)}^{-1.6}\big{(}\frac{\omega}{\omega_{J}}\big{)}^{0.8}\big{(}\frac{M_{p}}{M_{J}}\big{)}^{1.33}~{},$ (2)
where $a$ is the semi-major axis of the planet’s orbit in au, $\dot{M}$ ion is the stellar ionized mass-loss rate, $v_{\infty}$ is the terminal velocity of the stellar wind, and $\omega$ is the corotation speed (Farrell et al., 1999).
Terminal velocity of stellar wind can be calculated by
$v_{\infty}=0.75\times 617.5\times\sqrt{R_{\star}/M_{\star}}~{},$ (3)
where $R_{\star}$ and $M_{\star}$ are the host star radius and mass (O’Gorman et al., 2018). Therefore, the radio flux density ($S_{\nu}$) emitted by an exoplanet can be estimated using the following equation
$S_{\nu}=\frac{P_{rad}}{\Delta\nu\Omega d^{2}}~{},$ (4)
where $P_{rad}$ is the total radio emission power of the exoplanet (2), $d$ is the distance between the exoplanet and the observer, and $\nu$ is the frequency of the radio emission,
$\nu_{c}=23.5~{}\text{MHz}~{}\Big{(}\frac{\omega}{\omega_{J}}\Big{)}\Big{(}\frac{M_{p}}{M_{J}}\Big{)}^{5/3}\Big{(}\frac{R_{p}}{R_{J}}\Big{)}^{3}~{},$ (5)
where $R_{p}$ is the planetary radius. Substituting (4) into (2) gives
$S_{\nu}$ $=$ $7.6~{}\text{mJy}~{}(\frac{\omega}{\omega_{J}})^{-0.2}(\frac{M_{p}}{M_{J}})^{-0.33}(\frac{R_{p}}{R_{J}})^{-3}(\frac{\Omega}{1.6~{}sr})^{-1}(\frac{d}{10~{}pc})^{-2}$ (6)
$\times$ $(\frac{a}{1~{}au})^{-1.6}(\frac{{\dot{M}_{ion}}}{10^{-11}~{}M_{\odot}~{}yr^{-1}})^{0.8}(\frac{v_{\infty}}{100~{}km~{}s^{-1}})^{2}~{},$ (7)
where $\omega_{J}$, $M_{J}$, and $R_{J}$, are Jupiter’s corotation speed, mass, and radius and $\Omega$ is the beaming solid angle of the emission. In deriving this expression, we have followed (Farrell et al., 1999) and assumed that the planet will emit ECMI emission between the frequencies $0.3\nu_{c}$ and $\nu_{c}$, where $\nu_{c}$ is the maximum radiation frequency.
The planetary corotation speed ($\omega$) is a fundamental parameter in planetary magnetospheric dynamics. In the context of close-in exoplanets, where tidal locking is prevalent due to intense gravitational forces from the host star, the corotation rate becomes particularly significant. Tidal locking occurs when the orbital period of the planet matches its rotational period, resulting in one hemisphere of the planet permanently facing the star. As a consequence, the corotation rate of tidally locked planets is synchronized with their orbital motion. This synchronization affects the generation and distribution of electromagnetic phenomena within the magnetosphere, such as auroral emissions and field-aligned currents (Zarka et al., 2001; Seager & Hui, 2002). To account for tidal locking’s influence, we presume that exoplanets within orbits less than $0.1$ au are tidally locked, meaning their orbital period matches their corotation period. For exoplanets beyond this threshold, we employ the Darwin-Radau relation (Murray & Dermott, 2000)
$\omega=\sqrt{\frac{f~{}G~{}M_{p}}{R_{eq}^{2}}\Bigg{[}\frac{5}{2}(1-1.5~{}C)^{2}+\frac{2}{5}\Bigg{]}}$ (8)
where $C$ is 0.4 for rocky planets and for gas giant planets $C=0.25$ and $f$ is the planet’s oblateness (Hubbard, 1984). For the Jovian planet with orbital distance $0.1>a$ au, we use the oblateness of Jupiter ($f=0.064$), and for rocky planets, we use the Earth’s value which is $f=0.00335$ (Barnes & Fortney, 2003).
2.2 Detectablity of the Planetary Radio Signal
The leading contemporary ground-based telescope for detecting planetary radio emissions is the SKA telescope. Comprising two arrays, SKA1-Low and SKA1-Mid, it covers a broad frequency range from 50 MHz to 14 GHz. With its immense collecting area, totaling one square kilometer when fully operational, the SKA can capture faint radio signals from various cosmic phenomena, including nearby planets and distant galaxies. Its innovative design incorporates advanced signal processing techniques and data analysis algorithms to maximize scientific output while minimizing data volumes and processing requirements. As a collaborative effort involving over 20 countries, the SKA embodies international cooperation and scientific excellence, poised to unveil new frontiers in radio astronomy. For our observations using the SKA, we focus on the SKA1-Low and SKA1-Mid frequency ranges, up to 890 MHz. The imaging sensitivity for SKA1-Low and SKA1-Mid up to 890 MHz is detailed in Table LABEL:tab:SKA, considering a continuum observation with a fractional bandwidth of approximately $\Delta\nu/\nu_{c}\approx 0.3$, alongside an integration time of 1 hour.
Table 1: Image sensitivity of SKA1-Low and SKA1-mid within the indicated frequency bands for continuum observations (Braun et al., 2019)
$\nu_{min}$ [MHz] $\nu_{mid}$ [MHz] $\nu_{max}$ [MHz] $\sigma$ [$\mu Jy$]
50 60 69 163
69 82 96 47
96 114 132 26
132 158 183 18
183 218 253 14
253 302 350 11
350 410 480 16.8
480 560 650 8.1
650 770 890 4.4
We use the SKA image sensitivity for different ranges of frequencies as summarized in LABEL:tab:SKA.
3 Results
We use the NASA Exoplanet Archive to calculate the radio flux of the confirmed exoplanets. The 80 exoplanets have essential data such as planetary radius and mass, orbital distance, host star radius and mass, and the system’s distance from us information. To estimate $\dot{M}_{\text{ion}}$, we utilized typical mass loss ratios associated with the spectral types of the host stars (Michaud et al., 2011; Puls et al., 1996; Krtička, 2014; Wood et al., 2001; Wargelin & Drake, 2002). According to our analysis, Exoplanets Qatar-4 b, TOI-1278 b, and WASP-173 A b exhibit radio emissions that are detectable by the SKA telescope. All these three exoplanets are considered as close-in exoplanets. TOI-1278 b is a brown dwarf with an M-type host star. Its mass is 18.5 times that of Jupiter with an orbital period of approximately 14.5 days (0.095 au) (Artigau et al., 2021). WASP-173 A b is a hot Jupiter ($M_{p}=3.69\times~{}M_{J}$) orbits a G-type star with an orbital period 1.4 days. (Hellier et al., 2019). Qatar-4 b is also a hot Jupiter exoplanet ($M_{p}=6.1\times~{}M_{J}$) that orbits a K-type star with an orbital period of 1.8 days (Alsubai et al., 2017)
Figure 1 displays the distributions of planetary parameters, with exoplanets exhibiting detected radio fluxes depicted in distinct colors. It’s noticeable that a significant portion of exoplanets emitting intense radio fluxes do so at frequencies below the sensitivity range of the SKA telescope (below 50 MHz). One method to expand the detection of exoplanets with observable radio signals using current technology involves leveraging microlensing events. These events act as natural amplifiers, their magnification unaffected by the radiation’s wavelength. A notable characteristic of gravitational lensing is the preservation of radiation polarization during the event. Unlike unpolarized stellar plasma radiation, radio emissions from exoplanets exhibit circular polarization (Zarka et al., 2014). Consequently, microlensing events offer a potential means to enhance the detection of exoplanets with observable radio signals (Bagheri et al., 2024).
Figure 1: Characteristics of 80 confirmed exoplanets. Exoplanets exhibiting detectable radio signals are highlighted in different colors.
4 Discussion
Our understanding of planetary magnetic fields is predominantly derived from observations within our solar system. Diversifying our investigations and insights into exoplanetary magnetic fields could offer valuable perspectives on their internal dynamics and compositions.
In this study, employing the RBL model, we have identified three exoplanets—Qatar-4 b, TOI-1278 b, and WASP-173 A b—as potential candidates for detecting radio signals using the forthcoming SKA telescope. These exoplanets are characterized as hot Jupiters. Although ECMI can be generated in hot Jupiter magnetospheres, the escape of these radio emissions from these exoplanets may be hindered by their ionospheres (Koskinen et al., 2013; Weber et al., 2017). The generation of ECMI requires a specific balance between plasma and cyclotron frequencies, with a depleted and highly magnetized plasma being essential for emission (Grießmeier et al., 2007; Zarka et al., 2018). Close-in exoplanets experience significant atmospheric expansion due to factors like tidal forces and XUV-driven heating, leading to the formation of gaseous envelopes around the planets (Shaikhislamov et al., 2020; Johnstone et al., 2018, 2019). Studies suggest that the high plasma densities in these expanded atmospheres may prevent the escape of radio emissions or inhibit the generation of radio waves via ECMI (Weber et al., 2017). Therefore, detecting radio emissions from even one hot Jupiter would offer valuable insights into planetary responses to their environments and evolutionary paths, shedding light on the intricate interplay between magnetic fields, stellar wind, and atmospheric escape mechanisms. This knowledge would contribute significantly to our understanding of planet formation, migration, and potential habitability beyond our solar system.
Understanding the magnetic fields and magnetospheric emissions of exoplanets represents a cutting-edge scientific frontier for the coming decade, as emphasized in the Origins, Worlds, and Life Planetary Science & Astrobiology Decadal Survey report. This report identifies two Priority Science Question Topics that encompass aspects of planetary magnetic fields and their interaction with the solar wind. These topics are labeled as Q6: ”Solid Body Atmospheres, Exospheres, Magnetospheres, and Climate Evolution” and Q12.7: ”Exoplanets, Giant Planet Structure and Eevolution” (National Academies of Sciences et al., 2022)..
5 acknowledgements
We acknowledge the support of the US National Science Foundation (NSF) under Grant No. 2138122.